OpenStudy (anonymous):

I need help quickly Terrell's mom paid $27 for 18 juice boxes for the party. At that rate, how much would she pay for 24 juice boxes? A.$42 B.$36 C.$45 D.$51

OpenStudy (anonymous):

You want to set up a proportion. (27/18)=(x/24)

OpenStudy (anonymous):

Or you can think of it as canceling units ($27/18 boxes)*24 boxes the juice boxes cancel and you are left with the dollar amount

OpenStudy (anonymous):

so, it would be 27/18= 1.5 1.5/24= .0625 that is not one if the answers:(

OpenStudy (anonymous):

No, (27/18) = 1.5 1.5*24= Do you know how to cross multiply?

OpenStudy (anonymous):

not the best but yes

OpenStudy (anonymous):

if you were to use the proportion you would have 27*24=18x.

OpenStudy (anonymous):

oh,ok

OpenStudy (anonymous):

or 648=18x, then divide by 18 and you have x, the dollar amount

OpenStudy (anonymous):

i am not getting that sorry

OpenStudy (anonymous):

648/18 = x

OpenStudy (anonymous):

X=36

OpenStudy (anonymous):

yes, if each box costs 1.50 and you have 24 boxes its 36 dollars

OpenStudy (anonymous):

ooooooooooohhhhhhh, now i get this thanks